OpenStudy (anonymous):

identify the asymptotes of each function. then graph the function. f(x)=-1/x

OpenStudy (anonymous):

@welshfella @rational @texaschic101 @Abhisar @sammixboo @Donblue @Hero @KendrickLamar2014 @xo.minnie.xox @Nnesha @nechirwan @mathstudent55 @mathmate

OpenStudy (mathmate):

Hint: A vertical asymptote occurs when the denominator of a rational function becomes zero.

OpenStudy (anonymous):

ok so 0 because its a negative

OpenStudy (anonymous):

@mathmate

OpenStudy (mathmate):

Negative or not, a zero in the denominator makes x=0 an asymptote. The negative sign will affect the shape of the graph as x approaches zero from the negative side, and from the positive side. To give you an idea, I will sketch the graph of y=-1/(x-1) which has an asymptote at x=1 (i.e. x-1=0)